The length of a rectangle field is twice the width and the area is 1800 meters squared. Find the dimensions.
Can you please help me with this question? Thanks so much in advance:)

Answer by sofiyac(983)

The length of a rectangle field is twice the width and the area is 1800 meters squared. Find the dimensions.
**
let W=width
2W=length
Area=length*width
W*2W=1800

2W^2=1800
W^2=900
W=±√900
W=30
2W=60
width=30 m
length=60 m